OpenStudy (anonymous):

Simple Question: Simplify: x^3+8 I know the answer, so please just show steps so I can learn.. I forgot how to do this :/

OpenStudy (anonymous):

That looks like it in in simplest form to me.. unless it want's you to expand the x's?

OpenStudy (anonymous):

Yeah, I need it to look like (x+2) times something

OpenStudy (precal):

is it equal to something? did you write it incorrectly?

OpenStudy (anonymous):

Oh umm, I'm solving for a limit, limx->-2 (x+2)/(x^3+8)

OpenStudy (anonymous):

You need to complete the cube, using a formula that I forget

OpenStudy (anonymous):

So but I don't really understand how to get from x^3+8 to (x+2)(x^2-2x+4)

OpenStudy (anonymous):

(x+a)^3 = x^3 +3x^2 a + 3xa^2+a^3

OpenStudy (accessdenied):

A formula: \(a^3 + b^3 = (a + b)(a^2 - ab + b^2) \)

OpenStudy (anonymous):

his second number isn't a cube though, so that one won't help.

OpenStudy (accessdenied):

x^3 + 8 8 = 2^3 ?

OpenStudy (anonymous):

@AccessDenied It doesn't work that that don't forget that we are dealing with a polynominal

OpenStudy (precal):

how about just looking at the graph

OpenStudy (alexwee123):

factor (x^3+8) out :D

OpenStudy (anonymous):

Ohhh, thanks everyone I get it now. I used AccessDenied's formula :D

OpenStudy (anonymous):

oh right, 8=2^3.

OpenStudy (accessdenied):

a^3 + b^3 = (a + b)(a^2 - ab + b^2) a^2(a + b) - ab(a + b) + b^2 (a + b) a^3 + a^2 b - a^2 b + ab^2 + ab^2 + b^3 a^3 + b^3

OpenStudy (alexwee123):